# Reasoning Limitations of Multimodal Large Language Models. A case study of Bongard Problems

### v1

@@ 5 Experiments
Binary classification (Ground-truth). We first assess whether MLLMs can determine if a given concept matches a problem instance. On synthetic BPs, $3$ proprietary (GPT-4o, Gemini 1.5 Pro, Claude 3.5 Sonnet) and $2$ open-access (LLaVa-1.6 Mistral-7B, Pixtral 12B) models outperform a random classifier by a notable margin. On Bongard HOI, $3$ proprietary (GPT-4o, GPT-4 Turbo, Gemini 1.5 Pro) and the same $2$ open-access models also surpass random guessing. Notably, Pixtral 12B attained a perfect score on this dataset. On Bongard-OpenWorld GPT-4o, Gemini 1.5 Pro, LLaVa-1.6 Mistral-7B, and Pixtral 12B achieve reasonable results. Again, the leading model is Pixtral 12B with the outstanding $99\%$ outcome. Model accuracy drops significantly on Bongard-RWR, where only LLaVa-1.6 Mistral-7B and Pixtral 12B outperform a random classifier. This suggests that correctly identifying concepts expressed in Bongard-RWR likely requires more advanced reasoning abilities, even in the relatively simpler binary classification setting.
@@
Generative capabilities in the Direct setting. Model performance using the Direct generation strategy is generally weak on synthetic BPs, with the best model, GPT-4o, solving only 17 out of 100 problems. This indicates that the models struggle to identify abstract, synthetic concepts and express them in natural language. The challenge is even more apparent on Bongard-RWR, where the best model, GPT-4o, solves only 5 out of 60 problems. Nevertheless, performance improves on Bongard HOI and Bongard-OpenWorld, with best results of 35/100 and 40/100, resp. Notably, while GPT-4o achieves the highest scores on these two datasets, Pixtral 12B ranks second (28/100 and 33/100, resp.), showing that smaller open-access models can still be competitive in this setting. While the better performance on Bongard HOI and Bongard-OpenWorld may be attributed to a higher ratio of real-world images in the training data, the weak results on Bongard-RWR suggest that the discrepancy is more related to the specific underlying concepts than the visual domain as such (see Figure 7 in Appendix C for the details).
Independent image description. In the next experiment we analyse whether an iterative reasoning approach, in which the model first generates separate captions for each image and then combines them into a final answer, can improve performance. In comparison to the Direct strategy, the Descriptive one improves the best results across all datasets: from 17 to 21 on synthetic BPs, from 35 to 45 on Bongard HOI, from 40 to 57 on Bongard-OpenWorld, and from 5 to 13 on Bongard-RWR. A clear improvement can be observed individually for each proprietary model. The gain is less pronounced (if at all) for individual open-access models. We hypothesize that this improvement is due to the Descriptive setting being more aligned with the model’s training, where it primarily learns to caption individual images. However, this strategy doesn’t leverage the additional information present in the joint BP image, and certain context-dependent visual features may be missed in captioning. We believe that with further advancements in reasoning over multi-part compositional images, models in the Direct setting should eventually outperform the Descriptive strategy.
Contrastive reasoning. Correct identification of concepts in BPs requires a joint processing of the images from both sides of the problem. The Contrastive strategy evaluates the model ability to extract underlying differentiating concepts within such image pairs. Across all datasets, the models evaluated under the Contrastive strategy perform worse than with the Descriptive strategy. This points to the fundamental difference between human and machine approaches to solving AVR tasks. Humans often rely on direct comparisons between image panels from different categories to highlight differences (Nüssli et al., 2009), whereas the tested methods perform better when making comparisons on text-based image descriptions, potentially disregarding critical visual details missed during image captioning. This discrepancy indicates the need for further modeling improvements to fully leverage the Contrastive strategy.
Iterative reasoning. Next, we tested whether preserving responses from past turns in the dialog context could improve concept identification in both Descriptive and Contrastive settings. The Descriptive-Iterative strategy visibly worsens the results compared to its non-iterative counterpart across all datasets and models, except for negligible improvement of InternVL2-8B on synthetic BPs and several cases of a complete failure (accuracy of 00) for both strategies. In contrast, Contrastive-Iterative brings no improvement over Contrastive strategy in only $5$ cases, $2$ for synthetic BPs, and $3$ regarding Bongard-OpenWorld. Despite these improvements, Contrastive-iterative strategy generally performs worse than Descriptive. This indicates that contemporary models have difficulties to effectively use additional information from the context window.
Multimodal answer generation. In the final experiment, we assessed whether incorporating an image of the entire matrix at the answer generation step would improve the performance of Descriptive and Contrastive strategies. Descriptive-Direct shows performance improvements over Descriptive in $12$ out of $32$ (dataset, model) cases. Contrastive-Direct improves upon Contrastive in all (dataset, proprietary model) configurations, and additionally improves in certain (dataset, open-access model) settings. However, despite these gains, Contrastive-Direct overall performs worse than Descriptive, except for GPT-4o and InternVL2-8B on synthetic BPs, and InternVL2-8B on Bongard HOI. This suggests that contemporary models are to some extent capable of utilizing additional visual inputs to improve reasoning performance, in particular the newest GPT-4o, which displays improvement from incorporating the -direct extension in 7 out of 8 cases. Nevertheless, further work is needed to improve consistency across all models.
@@
Omni Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iter. 9 $7$ $4$ $\underline{8}$ $1$ 00 $1$ $2$
@@
Contrastive-iter. 20 $9$ $\underline{12}$ $11$ $1$ 00 00 $4$
@@
Omni Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iter. 32 $\underline{23}$ $6$ $6$ 00 $1$ $2$ $11$
@@
Contrastive-iter. 22 $20$ $15$ $\underline{21}$ $2$ $1$ $3$ $14$
@@
Omni Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iter. 31 $\underline{24}$ $6$ $13$ $4$ $2$ $2$ $11$
@@
Contrastive-iter. $\underline{25}$ $21$ $8$ 34 $5$ $1$ $3$ $18$
@@
Omni Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iter. 4 $1$ 00 $\underline{3}$ 00 00 00 00
@@
Contrastive-iter. 5 $3$ $3$ $\underline{4}$ 00 00 00 $1$
@@
Comparison of prompting strategies. Across all models, the Descriptive strategy achieves the highest scores on Bongard-RWR and Bongard-OpenWorld. In Bongard HOI, it ties with Descriptive-direct, while in synthetic BPs, it ranks just behind its -direct extension. As shown in Appendix D, altogether Descriptive strategies solve the same number of synthetic BPs as Contrastive strategies (44; Fig. 9), but lead in Bongard HOI (82 vs. 63; Fig. 11), Bongard-OpenWorld (90 vs. 76; Fig. 13), and Bongard-RWR (20 vs. 11; Fig. 15). This overall advantage of Descriptive over Contrastive strategies indicates that current MLLMs perform better with prompting strategies focused on processing single images. This also highlights the need to improve multi-image reasoning capabilities of MLLMs for tasks that require reasoning across multiple images. Figs. 9 – 16 in Appendix D further show that altogether the considered approaches solved 54, 89, 93, and 23 problems from synthetic BPs, Bongard HOI, Bongard-OpenWorld, and Bongard-RWR, resp. This raises the question of whether an ensemble combining all proposed strategies could further enhance model reasoning performance. We leave the exploration of this emerging direction for future work.
Proprietary vs. open-access models. Proprietary models generally outperform open-access ones, leading in 35 out of 40 (dataset, strategy) pairs. The black-box nature of proprietary models makes it challenging to attribute their advantage to specific aspects, whether it be the number of parameters, the size and composition of training data, or the pre- and post-processing methods. However, the recently released Pixtral 12B model performs competitively in multiple settings, occasionally surpassing proprietary models. This highlights the viability of developing competitive MLLMs without sacrificing accessibility. At the same time, a clear performance drop of Pixtral 12B on synthetic BPs and Bongard-RWR instances suggests its intrinsic weakness in reasoning about abstract concepts, whether reflected in synthetic or real-world manner, similarly to other open models.
Comparison with state-of-the-art. A direct comparison with the results from (Wu et al., 2024) is challenging due to the different ranges of test problems used in each study. With this caveat, we concentrate on key high-level observations from both works. Wu et al. (2024) primarily focus on a binary classification setting corresponding to the Images to Sides setup in our work. On Bongard-OpenWorld, our best performing models, Gemini 1.5 Pro and Claude 3.5 Sonnet, achieved 96% accuracy, while their top method, SNAIL—a meta-learning approach leveraging pre-trained OpenCLIP image representations—achieved 64%. This suggests that MLLMs, which uniformly process images and text, outperform decoupled two-stage approaches, which handle image captioning and text-based reasoning with different models. They also briefly consider a natural language generation task, where models describe concepts presented in the BP instance (Wu et al., 2024, Appendix E). They again use a two-step approach comprising fine-tuned BLIP-2 for image captioning and ChatGPT for concept generation. In contrast, we employ a single MLLM for both tasks. For evaluating free-form concept generation, they use automated text-based metrics, which provide a general measure of text similarity. We, however, employ a voting MLLM ensemble, offering a more direct assessment of solution correctness.
6 Conclusions
In this paper, we investigate the reasoning capabilities of proprietary and open-access MLLMs using BPs as a case study. Despite rapid progress, MLLMs still exhibit significant reasoning limitations. Across all proposed answer generation strategies, the best-performing model managed to solve only 22 out of 100 synthetic BPs. Performance moderately improves when dealing with real-world concepts, as demonstrate the results on the Bongard HOI and Bongard-OpenWorld datasets. However, the models still struggle to leverage additional multimodal data and effectively utilize the dialog context window. To delve deeper into the performance discrepancies between synthetic and real-world domains, we introduced Bongard-RWR, a new BP dataset designed to represent concepts from synthetic BPs via real-world images. Focused experiments with this dataset suggest that the models’ weak performance on synthetic BPs is not domain-specific but rather indicative of broader limitations in their reasoning abilities. On a positive note, experiments conducted in three binary classification settings show that some models achieve encouraging results, suggesting that current limitations may be overcome with future advancements.
Appendix A MLLM prompts
A.1 Prompts for Bongard-RWR generation
Prompt A.1 was used to select those images that correctly represent given concept translation. In addition to the left and right concepts, we also provided prompts briefly explaining the context that the image should match. These prompts were generated during the translation stage of our algorithm (see the fourth line in Algorithm 1).
Prompt 2: Prompt used for the selection of proper images for a translated concept. You translated a concept comparison from geometric domain to the real-world domain as follows: Geometric domain: <left_geometric_concept> vs <right_geometric_concept> Real world domain: {     "left": {         "concept": <left_concept>,         "prompt": <left_concept_description>     },     "right": {         "concept": <right_concept>,         "prompt": <right_concept_description>     } } Now, you need to check if the queried image matches your translation and provides enough information to distinguish it from the other concept. Don’t focus too much on the prompt. It’s just a hint for you to understand the concept better. Provided image represents <side_concept> Give your answer in the following format: EVALUATION: OK EXPLANATION: <here you can provide additional information> or EVALUATION: REJECTED EXPLANATION: <here you can provide additional information>
A.2 Prompt describing the Bongard problem
Prompt A.2 describing the BP task has been placed at the beginning of each solving strategy introduced in Section 3.1.
Prompt 3: Prompt explaining Bongard problems to an MLLM. A Bongard Problem is composed of left and right sides separated by a line. Each side contains six images. All images belonging to one side present a common concept, which is lacking in all images from the other side. The goal is to describe the rule that fits all images on the left side, but none on the right, and, conversely, the rule that fits all images on the right side, but none on the left. The description of the rule should be simple and concise. Example 1: All shapes on left are small. All shapes on right are big. Example 2: The left side contains circles. The right side contains triangles.
A.3 Prompts for Classification Strategies
Prompt A.3 was used to assess solution correctness (see Section 3.2). Prompt A.3 was used to assign images to sides (see Section 3.3).
Prompt 4: Prompt used for images to side classification (see Figure 3c). Two examples were provided to not bias the results of the model. You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to classify two test images to the corresponding side of the Bongard Problem, LEFT or RIGHT. Each image belongs to exactly one class. The test images belong to different classes. The images are always provided correctly. Respond only to the specific request. Respond in json using the following format. FIRST EXAMPLE: Left images: <small shapes> Right images: <big shapes> First test image: <small shape> Second test image: <big shape> Response: {     "first": {         "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.",         "concept": "small vs big",         "answer": "LEFT"     },     "second": {         "explanation": "The test image shows a big shape, similarly as all images on right. The images on left, on the other hand, feature small shapes.",         "concept": "small vs big",         "answer": "RIGHT"     } } END OF FIRST EXAMPLE SECOND EXAMPLE: Left images: <circles> Right images: <triangles> First test image: <triangle> Second test image: <circle> Response: {     "first": {         "explanation": "The test image shows a triangle, which matches all images on right. In contrast, the left side images feature circles.",         "concept": "circles vs triangles",         "answer": "RIGHT"     },     "second": {         "explanation": "The test image shows a circle, which matches all images on left. Conversely, the right side images feature triangles.",         "concept": "circles vs triangles",         "answer": "LEFT"     } } END OF SECOND EXAMPLE
Prompt 5: Prompt used for the solution correctness assessment (see Figure 3b). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to evaluate the correctness of the provided answer to the given Bongard Problem. All images are provided correctly. Do not explain the answer, just evaluate it. Respond ’OK’ if the answer is correct, otherwise respond ’WRONG’. User answer: <user_answer>
A.4 Prompts for Natural Language Answer Generation Strategies
Prompts A.4A.4 were used for natural language answer generation (see Section 3.1).
Prompt 6: Prompt used for the Direct strategy. (see Figure 2a). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem. What is the difference between the two sides of the problem?
Prompt 7: Prompt used to obtain the image descriptions in the Descriptive strategy (see Figure 2b). The provided image is a part of an abstract visual reasoning problem. Describe all crucial properties of the image. Your description should be as concise as possible. Focus on the most important details. The image is provided correctly. Respond only with descriptions.
Prompt 8: Prompt used for the Descriptive and Descriptive-direct strategies (see Figure 2b). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem using descriptions of its images. LEFT IMAGES: <left_descriptions> RIGHT IMAGES: <right_descriptions> What is the difference between the two sides of the problem?
Prompt 9: Prompt used to obtain the image descriptions in the Descriptive-iterative strategy (see Figure 2c). After the last image, we used the prompt: “That was the last image. Now provide your final answer.” You’ll receive a sequence of images that are a part of a single side of a Bongard Problem. The images will be provided one by one. Your goal is to find a common concept presented in all images. Your description should be as concise as possible. Focus on the most important details. Try to enhance the description of the concept after each image. The image is always provided correctly. Respond only to the specific request. The first image will be provided in the next message.
Prompt 10: Prompt used for the Descriptive-iterative strategy (see Figure 2c). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem using descriptions of two sides of the problem. LEFT SIDE DESCRIPTION: <left_description> RIGHT SIDE DESCRIPTION: <right_description> What is the difference between the two sides of the problem?
Prompt 11: Prompt used to obtain the comparison between the left and right image in the Contrastive strategy (see Figure 2d). After the last image, we used the prompt: “That was the last image. Now provide your final answer.” You are given two images extracted from the left and right side of a Bongard Problem, respectively. Your goal is to compare the images. Your comparison should be as concise as possible.
Prompt 12: Prompt used for the Contrastive and Contrastive-direct strategies (see Figure 2d). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem using comparisons between pairs of images. Each pair contains one image from the left and one from the right side of the problem. COMPARISONS: <comparisons> What is the difference between the two sides of the problem?
Prompt 13: Prompt used for the Contrastive-iterative strategy (see Figure 2e). After the last pair of images, we used the prompt: “It was the last pair of images. What is the difference between the two sides of the problem?” You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem. You’ll receive a sequence of image pairs. Each pair contains one image from the left and one from the right side of the problem. In each step compare the two images and refine the definitions of concepts that describe left and right sides of the problem. Your description should be as concise as possible. Focus on the most important details. The first pair will be provided in the next message.
Appendix B Evaluation of MLLMs answers

### v2

@@ Appendix C Extended Results
Binary classification (Ground-Truth). We assessed whether MLLMs can determine if a given concept matches a problem instance. On synthetic BPs, $3$ proprietary (GPT-4o, Gemini 1.5 Pro, Claude 3.5 Sonnet) and $2$ open-access (LLaVA-1.6 Mistral-7B, Pixtral 12B) models outperform a random classifier by a notable margin. On Bongard HOI, $3$ proprietary (GPT-4o, GPT-4 Turbo, Gemini 1.5 Pro) and the same $2$ open-access models also surpass random guessing. Notably, Pixtral 12B attained a perfect score on this dataset. On Bongard-OpenWorld GPT-4o, Gemini 1.5 Pro, LLaVA-1.6 Mistral-7B, and Pixtral 12B achieve reasonable results. Again, the leading model is Pixtral 12B with the outstanding $99\%$ outcome. Model accuracy drops significantly on Bongard-RWR, where only LLaVA-1.6 Mistral-7B and Pixtral 12B outperform a random classifier. This suggests that correctly identifying concepts expressed in Bongard-RWR likely requires more advanced reasoning abilities, even in the relatively simpler binary classification setting.
@@
Binary classification on RWR variants. In the Ground-Truth and Incorrect Label settings, models performed similarly across the Bongard-RWR dataset and its three variants (RWR-G, RWR-S, RWR-SG), indicating that color removal and cropping did not substantially affect performance in these settings. However, in the Images to Sides setup (Table 2), several models showed clear differences. For example, Phi-3.5-Vision improved from solving $1$ problem on RWR to $16$ on RWR-G, $8$ on RWR-S, and $18$ on RWR-SG, suggesting that color removal and cropping reduce unnecessary variation, and shift the focus of the model toward structural patterns. Encouraged by these improvements, we evaluated larger model variants, including InternVL2.5 38B and 78B (Chen et al., 2024a, c), Qwen2VL 72B-Instruct (Bai et al., 2023; Wang et al., 2024), and LLaVA-NeXT 110B, which showed similar trends. Overall, the RWR-G variant highlighted models’ sensitivity to color, while RWR-S and RWR-SG revealed how cropping affects performance by reducing empty space between panels. Nevertheless, in most cases the models were unable to surpass the random guess threshold, highlighting their inherent limitations in recognizing Bongard-RWR concepts.
Figure 9: Concept selection. Multi-class classification accuracy of MLLMs across datasets and $k\in\{2,4,8,16\}$ choices. Detailed results are presented in Table 3.
Table 3: Concept selection. Multi-class classification accuracy of MLLMs across datasets and $k\in\{2,4,8,16\}$ choices. InternVL2 is denoted as IVL2, InternVL2.5 as IVL2.5, and Qwen2VL Instruct as Q2VL. High-level trends are illustrated in Fig. 9.
Bongard HOI B-OpenWorld Bongard-RWR Synthetic BPs
$2$ $4$ $8$ $16$ $2$ $4$ $8$ $16$ $2$ $4$ $8$ $16$ $2$ $4$ $8$ $16$
IVL2 8B $0.95$ $0.97$ $0.90$ $0.90$ $0.92$ $0.89$ $0.92$ $0.81$ $0.48$ $0.50$ $0.25$ $0.25$ $0.61$ $0.49$ $0.30$ $0.16$
IVL2.5 8B $0.97$ $0.96$ $0.92$ $0.93$ $0.96$ $0.96$ $0.94$ $0.87$ $0.82$ $0.52$ $0.45$ $0.27$ $0.73$ $0.56$ $0.41$ $0.30$
IVL2.5 38B $1.00$ $0.97$ $0.99$ $0.97$ $1.00$ $1.00$ $0.99$ $0.98$ $0.73$ $0.65$ $0.52$ $0.32$ $0.82$ $0.71$ $0.66$ $0.46$
IVL2.5 78B $1.00$ $0.99$ $0.98$ $0.99$ $0.99$ $1.00$ $0.99$ $1.00$ $0.88$ $0.65$ $0.68$ $0.38$ $0.86$ $0.83$ $0.75$ $0.70$
Q2VL 72B $1.00$ $0.99$ $0.95$ $0.96$ $0.98$ $0.98$ $0.99$ $0.97$ $0.85$ $0.67$ $0.57$ $0.32$ $0.91$ $0.69$ $0.59$ $0.41$
Multi-class classification (concept selection). Fig. 9 and Table 3 present detailed results of the multi-class classification experiments, where models were tasked with selecting the correct concept from $k\in\{2,4,8,16\}$ choices across all four datasets. First, the performance improves from InternVL2 8B to InternVL2.5 8B, reflecting improvements in training data and model architecture in newer model versions. Next, moving from InternVL2.5 8B to InternVL2.5 38B and InternVL2.5 78B shows further gains, highlighting the benefits of scaling up model size, which we explore further in Appendix D. Comparing InternVL2.5 78B with Qwen2VL 72B-Instruct reveals that state-of-the-art models also from other families face challenges on this task, particularly on Bongard-RWR and synthetic BPs. It’s worth noting that the average human accuracy of $65\%$ on Bongard-RWR observed in the free-form setup is similar to the performance of best models in the much simpler multiple-choice setup with only $k=4$ candidate answers. This observation underscores the limitations of current models compared to humans.
@@
o Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iterative 9 $7$ $4$ $\underline{8}$ $1$ 00 $1$ $2$
@@
Contrastive-iterative 20 $9$ $\underline{12}$ $11$ $1$ 00 00 $4$
@@
o Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iterative 32 $\underline{23}$ $6$ $6$ 00 $1$ $2$ $11$
@@
Contrastive-iterative 22 $20$ $15$ $\underline{21}$ $2$ $1$ $3$ $14$
@@
o Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iterative 31 $\underline{24}$ $6$ $13$ $4$ $2$ $2$ $11$
@@
Contrastive-iterative $\underline{25}$ $21$ $8$ 34 $5$ $1$ $3$ $18$
@@
o Turbo 1.5 Pro 3.5 Sonnet 8B Mistral 7B 3.5V 12B
@@
Descriptive-iterative 4 $1$ 00 $\underline{3}$ 00 00 00 00
@@
Contrastive-iterative 5 $3$ $3$ $\underline{4}$ 00 00 00 $1$
@@
Appendix D The Impact of Model Scaling on Abstract Reasoning Abilities
Figure 10: The impact of model scaling on abstract reasoning abilities. We considered a diverse set of model sizes across proprietary and open-access MLLMs. The experiments cover all $4$ datasets using the Direct and Descriptive solution strategies.
Performance of MLLMs on downstream tasks is often correlated with the number of model parameters and the size of training datasets (Kaplan et al., 2020; Hoffmann et al., 2022). To investigate the relationship between model scaling and abstract reasoning performance, we conducted experiments with a diverse set of model sizes across proprietary and open-access MLLMs. To this end, we evaluated both smaller and larger variants of the selected models. Specifically, we considered GPT-4o mini and Gemini 1.5 Flash as smaller counterparts to GPT-4o and Gemini 1.5 Pro, resp. Also, we tested multiple configurations of InternVL2 and LLaVA-NeXT model families including InternVL2-8B, InternVL2-26B, InternVL2-40B, InternVL2-Llama3-76B, LLaVA-v1.6 Vicuna-13B, LLaVA-v1.6 34B, LLaVA-NeXT 72B, and LLaVA-NeXT 110B. We conducted experiments on all $4$ datasets using two solution strategies, including Direct, which is an intuitive baseline, and Descriptive, the most effective strategy identified in the main experiments.
The results are presented in Fig. 10. In general, larger proprietary models outperformed their smaller counterparts in $10$ out of $16$ cases. However, smaller variants sometimes performed better than larger ones. For instance, on Bongard HOI with the Direct strategy, GPT-4o mini and Gemini 1.5 Flash surpassed their larger alternatives. This suggests that smaller models can achieve competitive performance in abstract reasoning.
For open-access models, performance consistently improved with model size. For example, the results of InternVL2 on Bongard HOI increased from $12$ to $25$ and from $2$ to $29$ for Direct and Descriptive strategies, resp. Similarly, on Bongard HOI, the performance of LLaVA-NeXT improved from $5$ to $27$ and from $4$ to $27$ for the two strategies. Analogous improvements were observed on Bongard-OpenWorld, highlighting the potential benefits of model scaling.
Despite these significant improvements in open-access models, proprietary models consistently outperformed them. In particular, GPT-4o mini achieved worse results than the best open-access model in a single case only, i.e., Bongard HOI using the Descriptive strategy ($26$ vs. $29$). Although model scaling demonstrates its potential to enhance abstract reasoning, as shown by the open-access models, the relatively strong performance of GPT-4o mini shows that a large parameter count is not necessarily critical for excelling in abstract reasoning tasks. Consequently, these results suggest that simply scaling model size may be insufficient to achieve stronger abstract reasoning capabilities and future efforts should explicitly address this aspect, e.g., by incorporating AVR datasets into model training.
Appendix E Evaluation of MLLMs Answers